OpenStudy (anonymous):

│x – 5│ = 3 a. x = 8 and x = 2 b. x = 8 and x = –2 c. x = 8 d. x = 2

OpenStudy (anonymous):

if |x-5| >0 , then x-5=3 if |x-5| <0, then -x-(-5)=3; -x+5=3

OpenStudy (anonymous):

since x is a variable, you can't say |x-5| is > or < than 0, so you must consider both cases (unless the question tells you if it is < or >)

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

not d

OpenStudy (anonymous):

ohh a?

OpenStudy (anonymous):

haha right. It's a

OpenStudy (anonymous):

linear function means the graph will be a straight line so, if you take the same deltax for diferent x's, there should be the same deltay in both cases I guess it's a bit confuse - sorry I can't explain the right way cause my english is poor...

OpenStudy (anonymous):

like if you pick up x1=1, y1=1 & x2=2, y2=-3 for any other deltax = x2-x1 = 1, there should be the same deltay = y2=y1 = -4

OpenStudy (anonymous):

i suppose its not ?

OpenStudy (anonymous):

it is! plot the graph and you will see. Every point (x,y) is in an unique line

OpenStudy (anonymous):

oooh haha thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

Tell whether the sequence is arithmetic. If it is, what is the common difference? –12, –7, –2, 3, . . . (1 point) yes; 5 yes; 8 yes; 3 no

OpenStudy (anonymous):

and you can even write the function, which is y=5-4x

OpenStudy (anonymous):

that one is yes; 5 ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. x y 3 6 6 18 8 24 (1 point) yes; y = 2x yes; y = 3x yes; y = 4x no; y does not vary directly with x

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

in this case, any member of the sequence is given by An= -12 + 5n

OpenStudy (anonymous):

hahaha, calm down!

OpenStudy (anonymous):

haha mybad

OpenStudy (anonymous):

well, that's the same as the other

OpenStudy (anonymous):

actually, it is yes

OpenStudy (anonymous):

The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time? Time (months) Height (inches) 2 14 4 23 6 32 8 41 y = 9x + 2 y = 9/2x + 5 y = 9/2x y = 5x + 9/2

OpenStudy (anonymous):

(by the way, the previous one is d, sorry)

OpenStudy (anonymous):

i know, thank you

OpenStudy (anonymous):

you need to find the slope which is deltay/deltax

OpenStudy (anonymous):

y=ax+b a is the slope once you find it, just put any value for y from the tableand you get b

OpenStudy (anonymous):

Write y = –2/5x – 3 in standard form using integers. 2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15

OpenStudy (anonymous):

it is (-2/5)x , right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

just multiply everything by 5

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

no a ! haha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks your helping so much

OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y = –1/3x + 5; (4, 3) y = –3x + 15 y = 3x – 9 y = –4x + 3 y = –1/3x – 11

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

aaaaaa, I forgot the rule

OpenStudy (anonymous):

to be perpendicular, the multiplication of the slopes' values must be -1, is it right?

OpenStudy (anonymous):

the equation you need to find is y=ax+b a.(-1/3) = ? then substitute (4;3)

OpenStudy (anonymous):

ah, the slope2 of a perpendicular line is -1/slope1

OpenStudy (anonymous):

a1 the slope of a line a2 the slope of a perpendicular line then a2= -1/a1 or a1.a2=-1

OpenStudy (anonymous):

hmm C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the solution of the system? y = x – 14 y = –4x + 11 (5, –9) 1 over 5 and – (69 over 5) 5 over 3 and – (37 over 3) – (3 over 5) and – (73 over 5)

OpenStudy (anonymous):

subtracting the first by the second, you have 0=5x-25 this gives you x=5

OpenStudy (anonymous):

substitue x in any of the equations and you find y

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

idk where 0=5x-25 is from :/

OpenStudy (anonymous):

y=x-14 y=-4x+11 (minus) _____________ 0=x-(-4x)-14-11

OpenStudy (anonymous):

or since y=y, then x-14=-4x+11

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

no, it's a, the simplest

OpenStudy (anonymous):

ohh lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Bradley and Ryan are having a contest to see who can collect the most bumper stickers. Bradley started with 23 bumper stickers and plans to add 4 each month. Ryan has 9 bumper stickers and plans to add 6 each month. After how many months will they have the same number of bumper stickers? (1 point) 10 months 2 months 7 months 51 months

OpenStudy (anonymous):

a way to solve it is writing the equations of stickers collected by each of them

OpenStudy (anonymous):

being m = months, then Bradley: Stickers=23+4m Ryan: Stickers=9+6m

OpenStudy (anonymous):

it means BRadley has 23 stickers and owns 4more each month and Ryan has 9 stickers and owns 6 more each month

OpenStudy (anonymous):

b.2 months?

OpenStudy (anonymous):

it is the same as the previous, and you can just write y=23+4x y=9+6x you want to know x

OpenStudy (anonymous):

not b

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

in this case, you don't want to know y

OpenStudy (anonymous):

either, you could substitute the options in the formulas and see if they give the same number

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

Sure you know! Bradley: [number of stickers]=23+4x Ryan: [number of stickers]=9+6x

OpenStudy (anonymous):

if you call number of stickers as 'y' then you see that y MUST be the same in both equations